OpenStudy (anonymous):

According to the theorem of convergence of the power series for what possible values of x can the series converge?:

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} a_{n}x^{n}\]

OpenStudy (anonymous):

For every power series, there is a number R \[ 0\le R \le \infty \] called the radius of convergence of the series. For \[|x| < R \] the power series converges absolutely. For |x| = R, the situation must be examined separately.